A Bibliography of Books about the Internet and Networking: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 August 2024 Version 1.20

Title word cross-reference

Accurate [LNNNC16]. agenda [LM15]. algorithmic [LPS11]. analysis [MST^{+}19]. application [Mar18]. Applied [Boy10]. approach [KR10]. assessments [Wat14]. attacks [PKB10]. auditing [Jac10].

band [Eva18]. basics [EB11, Eng13]. battles [Sch15]. before [Gre18]. benefits [Til12]. big [LSBN14]. Broad [Eva18]. broadband [LM15]. bug [Mar18]. Built [Car13].

center [Blu12]. certified [DeF10]. channels [MST^{+}19]. CISO [Til12]. Cleaning [IC19]. cloud [Gre18]. collect [Sch15]. Computer [Bas13, JI12, KR10, Mil11]. computing [Gre18]. control [Sch15]. Countermeasures [FG10]. creating [Wil10]. Cyber [CK10]. cybercrime [Kre14]. Cyberwar [Gre19].

Dangerous [Gre19]. Data [IC19, Sch15, LSBN14]. deadliest [PKB10]. Decade [SMZ^{+}10]. defense [Wat14]. devices [PR15]. Digital [SMZ^{+}10, JI12, LM15]. discover [Mar18]. do [CK10]. door [Kre14]. down [KR10]. Dynamics [LM15].

easy [EB11, Eng13]. engagement [LSBN14]. engineering [Wat14]. epidemic [Kre14]. Era [Gre19]. Essays [SMZ^{+}10]. Ethical [Bal14, DeF10, EB11, Eng13, Mar18]. Europe [LM15]. European [Mar12, Oli12]. executing [Wat14]. execution [MST^{+}19]. exposed [BSRS13].

Facets [FG10]. Financial [LNNNC16]. flaws [Mar18]. formal [Wil10]. foundations [LPS11]. Fragility [Tuf17]. framework [Til12]. frameworks [LSBN14]. front [Kre14]. Future [SMZ^{+}10].

Gas [Tuf17]. Geeks [Car13]. global [Kre14]. Goliath [Sch15]. good [LSBN14]. guide [Bal14, Boy10, DeF10, Mar18, Til12].

hacker [DeF10]. Hackers [Gre19, Mar18]. Hacking [BSRS13, BB14, PR15, Bal14, EB11, Eng13, Wei14, Wil10, WA11]. Hands [Mar18, Boy10, Wei14]. Hands-on [Mar18, Boy10, Wei14]. Happen [DB10]. help [Mar18]. here [MST^{+}19]. hidden [Mar12, Sch15, Oli12]. History [DB10]. Homeland [FG10]. Hunt [Gre19]. hunting [Mar18].

information [Boy10, Mjö12, Vac10]. International [DB10]. Internet [SMZ^{+}10, Blu12, Car13, Eva18, LPS11]. introduction [Mjö12, Wei14]. Issue [FG10].

journey [Blu12].

Kali [BB14]. Kremlin [Gre19].

lab [Wil10]. literacy [JI12]. low [PR15].

Made [DB10, EB11, Eng13, Eva18]. manage [Til12]. Managing [Vac10]. Markets [LNNNC16, LM15]. Mathematical [LPS11]. maximize [Til12]. mobile [BSRS13]. Most [Gre19]. multidisciplinary [Mjö12].

nation [Kre14]. national [CK10]. Network [Car13, Jac10, Mil11, PKB10, Mil11]. Networked [Tuf17]. Networking [DB10, Mar12, Oli12, KR10]. Next [SMZ^{+}10, CK10]. Ninja [WA11].

Official [DeF10]. operating [Wil10]. organized [Kre14]. origins [Gre18].

pen [Wat14]. Penetration [Ano11, Wei14, Bal14, Bas13, BB14, EB11, Eng13, Mar18, NHS13, PR15, Til12, Wat14, Wil10, WA11, Wil13]. People [DB10]. plan [Til12]. Power [Tuf17, PR15]. practical [BB14, Mar18]. Prehistory [Mar12, Oli12]. Privacy [FG10, LSBN14]. Professional [Wil10, Wil13]. Protest [Tuf17]. Protocol [Mil11]. public [LSBN14].

realizing [LM15]. reconnaissance [NHS13]. Research [DB10, Mar12, Oli12]. review [DeF10].

safe [JI12]. Sandworm [Gre19]. secrets [BSRS13]. Security [FG10, Bas13, BSRS13, Boy10, CK10, Jac10, JI12, Mar18, Mjö12, Vac10]. Seven [PKB10]. side [MST^{+}19]. side-channels [MST^{+}19]. Social [Wat14]. software [Boy10]. solutions [BSRS13]. Spam [Kre14]. Spectre [MST^{+}19]. speculative [MST^{+}19]. stay [MST^{+}19]. staying [JI12]. story [Eva18, Kre14]. Synchronization [LNNNC16, Mil11].

tactics [WA11]. Tear [Tuf17]. techniques [BB14, WA11]. testers [Mar18]. testing [Ano11, Bal14, Bas13, BB14, EB11, Eng13, NHS13, PR15, Til12, Wat14, Wei14, Wil10, WA11, Wil13]. tests [Wat14]. threat [CK10]. Threats [FG10]. Time [LNNNC16, Mil11]. top [KR10]. top-down [KR10]. Traceable [LNNNC16]. Tubes [Blu12]. Twitter [Tuf17]. Tym [Gre18].

unconventional [WA11]. untold [Eva18, Gre18].

Verifiable [LNNNC16].

war [CK10]. web [Mar18]. who [DB10, Eva18]. Wireless [NHS13]. women [Eva18]. World [LNNNC16, JI12, Sch15].

References

Anonymous:2011:PT

[Ano11] Anonymous. Penetration testing. Course Technology, Cengage Learning, Clifton Park, NY, USA, 2011. ISBN 1-4354-8366-9 (vol. 1 : paperback), 1-4354-8367-7 (vol. 2 : paperback), 1-4354-8368-5 (vol. 3 : paperback), 1-4354-8369-3 (vol. 4 : paperback), 1-4354-8370-7 (vol. 5 : paperback). ???? pp. LCCN QA76.9.A25 P4355 2011.

Baloch:2014:EHP

[Bal14] Rafay Baloch. Ethical hacking and penetration testing guide. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4822-3161-1 (paperback). ???? pp. LCCN QA76.9.A25 B356 2014.

Basta:2013:CSP

[Bas13] Alfred Basta. Computer security and penetration testing. Course Technology, Cengage Learning, Boston, MA, second edition, 2013. ISBN 0-8400-2093-7. ???? pp. LCCN ????

Broad:2014:HKP

[BB14] James Broad and Andrew Bindner. Hacking with Kali: practical penetration testing techniques. Syngress, an imprint of Elsevier, Amsterdam, The Netherlands, first edition, 2014. ISBN 0-12-407749-8 (alkaline paper). ix + 227 pp. LCCN QA76.9.A25 B772 2014.

Blum:2012:TJC

[Blu12] Andrew Blum. Tubes: a journey to the center of the Internet. Ecco, New York, NY, USA, 2012. ISBN 0-06-199493-6. 294 pp. LCCN TK5105.875.I57 B58 2012.

Boyle:2010:AIS

[Boy10]  Randall Boyle.  Applied in-

formation security: a handson guide to information security software. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2010. ISBN 0-13-612203-5. viii + 226 pp. LCCN QA76.9.A25.

Bergman:2013:HEM

[BSRS13] Neil Bergman, Mike Stanfield, Jason Rouse, and Joel Scambray. Hacking exposed: mobile security secrets and solutions. Mc-Graw-Hill, New York, NY, USA, 2013. ISBN 0-07-181701-8 (paperback). xxvii + 289 pp. LCCN TK5105.59 .B464 2013.

Carpenter:2013:NGH

[Car13] Brian E. Carpenter. Network Geeks — How They Built the Internet. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4471-5024-4 (paperback), 1-4471-5025-2 (ebook). ix + 161 + 15 pp. LCCN QA76.17; TK5105.875.I57 C37

2013. URL https://sites.
google.com/site/bcabrc/network-
geeks-book.

Clarke:2010:CWN

[CK10] Richard A. (Richard Alan)

Clarke and Robert K. Knake. Cyber war: the next threat to national security and what to do about it. Ecco, New York, NY, USA, 2010. ISBN 0-06-196223-6 (hardcover). xiv + 290 pp. LCCN HV6773.2 .C53 2010.

Davies:2010:HIR

[DB10] Howard Davies and Beatrice Bressan, editors. A History of International Research Networking: The People who Made it Happen. Wiley, New York, NY, USA, February 2010. ISBN 3-527-32710-X (hardcover), 3-527-62933-5 (e-book). xxviii + 317 pp. LCCN TK5105.875.I57 H57 2010.

DeFino:2010:OCE

[DeF10] Steven DeFino. Official certified ethical hacker review guide. Course Technology, Cengage Learning, Boston, MA, USA, 2010. ISBN 1-4354-8853-9 (paperback). xxii + 361 pp. LCCN QA76.3 .D445 2010.

Engebretson:2011:BHP

[EB11] Pat (Patrick Henry) Engebretson and James Broad, CISSP., editors. The basics of hacking and penetration testing: ethical hacking and penetration testing made easy. Syngress the basics. Syngress Publishing, Inc., Rockland, MA, USA, 2011. ISBN 1-59749-655-3. xvii + 159 pp. LCCN QA76.9.A25 E5443 2011.

Engebretson:2013:BHP

[Eng13] Pat (Patrick Henry) Engebretson, editor. The basics of hacking and penetration testing: ethical hacking and penetration testing made easy. Syngress, an imprint of Elsevier, Amsterdam, The Netherlands, second edition, 2013. ISBN 0-12-411644-2 (paperback). xviii + 204 pp. LCCN QA76.9.A25 E5443 2013.

Evans:2018:BBU

[Eva18] Claire Lisa Evans. Broad band: the untold story of the women who made the Internet. Portfolio, New York, NY, USA, 2018. ISBN 0-7352-1175-2 (hardcover), 0-7352-1176-0 (e-pub). 278 pp. LCCN QA76.2.A2 E93 2018.

Franceschetti:2010:HSF

[FG10] Giorgio Franceschetti and Marina Grossi, editors. Homeland Security Facets Threats: Countermeasures, and the Privacy Issue. Artech House Inc., Norwood, MA, USA, 2010. ISBN 1-60807-106-5. 280 (est.)

pp.  LCCN ????  URL

                                        http://www.artechhouse.com/
                                        Detail.aspx?strIsbn=978-1-
                                        60807-106-7.

Gregory:2018:TBUb

[Gre18] Nathan Gregory. The Tym before: the untold origins of cloud computing. lulu.com, Lexington, KY, USA, 2018. ISBN 1-973430-18-5. 476 pp. LCCN QA76.585 .G74 2018.

Greenberg:2019:SNE

[Gre19] Andy Greenberg. Sandworm:

a New Era of Cyberwar and the Hunt for the Kremlin’s Most Dangerous Hackers. Doubleday, New York, NY, USA, 2019. ISBN 0-385-54440-5 (hardcover), 0-525-56463-2 (paperback), 0-385-54441-3 (e-book). xiii + 348 pp. LCCN HV6773.R8 G74 2019.

Ilyas:2019:DCf

[IC19] Ihab F. Ilyas and Xu Chu. Data Cleaning, volume 28 of ACM Books. ACM Press, New York, NY 10036, USA, 2019. ISBN 1-4503-7152-3 (hardcover), 1-4503-7153-1 (paperback), 1-4503-7154-X (e-pub), 1-4503-7155-8 (ebook). xix + 260 pp. LCCN QA76.9.D3 I483 2019.

Jackson:2010:NSA

[Jac10] Chris Jackson. Network security auditing. Cisco Press networking technology series. Cisco Press, Indianapolis, IN, USA, 2010. ISBN 1-58705-352-7 (paperback). xxiv + 488 pp. LCCN TK5105.59 .J325 2010.

Jacobson:2012:CSL

[JI12] Douglas Jacobson and Joseph Idziorek. Computer security literacy: staying safe in a digital world. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-5618-4 (paperback). ???? pp. LCCN QA76.9.A25 J224

Kurose:2010:CNT

[KR10] James F. Kurose and Keith W. Ross. Computer networking: a top-down approach. Addison-Wesley, Reading, MA, USA, fifth edition, 2010. ISBN 0-13-607967-9. xxiv + 862 pp. LCCN TK5105.875.I57 K88 2010.

Krebs:2014:SNI

[Kre14] Brian Krebs. Spam nation: the inside story of organized cybercrime — from global epidemic to your front door. Sourcebooks, Inc., Naperville, IL, USA, 2014. ISBN 1-4022-9561-8 (hardcover). ???? pp. LCCN HV6773.2 .K74 2014.

Lemstra:2015:DBM

[LM15] W. (Wolter) Lemstra and William H. Melody, editors. The Dynamics of broadband markets in Europe: realizing the 2020 digital agenda. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-07358-8 (hardcover). xx + 409 pp. LCCN HE8084 .D96

                                        2015. URL http://assets.
                                        cambridge.org/97811070/73586/
                                        cover/9781107073586.jpg.

Lombardi:2016:ATV

[LNNNC16] Michael A. Lombardi, Andrew N. Novick, George Neville-Neil, and Ben Cooke. Accurate, traceable, and verifiable time synchronization for world financial markets. Journal of research of the National Institute of Standards and Technology, 121(??):

436–463, October 7, 2016. CO-DEN JRITEF. ISSN 1044-677X (print), 2165-7254 (elec-

tronic). URL http://nvlpubs.
nist.gov/nistpubs/jres/121/
jres.121.023.pdf.

Luccio:2011:MAF

[LPS11] Fabrizio Luccio, Linda Pagli, and Graham Steel. Mathematical and algorithmic foundations of the Internet. Chapman and Hall/CRC Press applied algorithms and data structures series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2011. ISBN 1-4398-3138-6 (hardback). xv + 205 pp. LCCN TK5105.875.I57 L835 2011.

Lane:2014:PBD

[LSBN14] Julia I. Lane, Victoria Stodden, Stefan Bender, and Helen Nissenbaum, editors. Privacy, big data, and the public good: frameworks for engagement. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-06735-9 (hardcover), 1-107-63768-6 (paperback). xix + 322 pp. LCCN JC596 .P747 2015.

Martin:2012:HPE

[Mar12] Olivier H. Martin. The “hidden”

prehistory of European research networking. Report, ICTC Consulting, ????, Switzerland, October 25, 2012. 125 pp. URL http:

//ictconsulting.ch/reports/
European-Research-Internet-
History.pdf.

Marshall:2018:HBH

[Mar18]  Joseph Marshall.  Hands-on

bug hunting for penetration testers: a practical guide to help ethical hackers discover web application security flaws. Packt Publishing, Birmingham, 2018. ISBN 1-78934-420-4 (paperback), 1-78934-989-3 (ebook). viii + 234 pp. LCCN TK5105.5485. URL https:

                                        //international.scholarvox.
                                        com/book/88863226.

Mills:2011:CNT

[Mil11] David L. Mills. Computer network time synchronization: the Network Time Protocol. Taylor and Francis, Boca Raton, FL, USA, second edition, 2011. ISBN 1-4398-1463-5. ???? pp. LCCN TK5105.575 .M55 2011.

Mjolsnes:2012:MII

[Mjö12] Stig F. Mjölsnes, editor. A multidisciplinary introduction to information security. Discrete mathematics and its applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4200-8590-5 (hardback). xxv + 322 pp. LCCN QA76.9.A25 M845 2012.

Mcilroy:2019:SHS

[MST^{+}19] Ross Mcilroy, Jaroslav Sevcik, Tobias Tebbi, Ben L. Titzer, and Toon Verwaest. Spectre is here to stay: An analysis of side-channels and speculative execution. arxiv.org, ??(??):1–26, February 14, 2019. URL https:

//arxiv.org/abs/1902.05178.

Neely:2013:WRP

[NHS13] Matthew Neely, Alex Hamerstone, and Chris Sanyk. Wireless reconnaissance in penetration testing. Elsevier/Syngress, Amsterdam, The Netherlands, 2013. ISBN 1-59749-731-2. xvi + 166 pp. LCCN TK7882.E2 N44 2013.

Olivier:2012:HPE

[Oli12]  Martin Olivier.  “Hidden”

Prehistory of European Research Networking. Trafford Publishing, [unknown], 2012. ISBN 1-4669-3872-2. LCCN

????  URL http://www.
ictconsulting.ch/reports/European-
Research-Internet-History.
pdf;   https://www.amazon.
com/Hidden-Prehistory-European-
Research-Networking-ebook/
dp/B0792XPN4G.

Prowell:2010:SDN

[PKB10] Stacy J. Prowell, Rob Kraus, and Mike Borkin. Seven deadliest network attacks. Syngress seven deadliest attacks series. Syngress Publishing, Inc., Rockland, MA, USA, 2010. ISBN 1-59749-549-2 (paperback). xiv + 142 pp. LCCN TK5105.59 .P76 2010.

Polstra:2015:HPT

[PR15] Philip Polstra and Vivek Ramachandran, editors. Hacking and penetration testing with low power devices. Elsevier, Amsterdam, The Netherlands, 2015. ISBN 0-12-800751-6, 0-12-800824-5 (e-book), 1-322-09763-1 (e-book). xv + 243 pp. LCCN QA76.9.A25 P5965 2015. URL http://www.

sciencedirect.com/science/
book/9780128007518.

Schneier:2015:DGH

[Sch15] Bruce Schneier. Data and Goliath: the hidden battles to collect your data and control your world. W. W. Norton & Co., New York, NY, USA, 2015. ISBN 0-393-24481-4 (hardcover). 383 pp. LCCN HM846 .S362 2015.

Szoka:2010:NDD

[SMZ^{+}10] Berin Szoka, Adam Marcus, Jonathan L. Zittrain, Yochai Benkler, and John G. Palfrey. The Next Digital Decade: Essays on the Future of the Internet. TechFreedom, Washington D.C, 2010. ISBN 1-4357-6786-1. 575 pp. LCCN HM851 .N49 2010.

URL https://books.google.
com/books?id=DZMSQHOfUKcC.

Tiller:2012:CGP

[Til12] James S. Tiller. CISO’s guide to penetration testing: a framework to plan, manage, and maximize benefits. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-8027-1 (hardcover). xiv + 374 pp. LCCN QA76.9.A25 T56 2012.

Tufekci:2017:TTG

[Tuf17] Zeynep Tufekci. Twitter and Tear Gas: the Power and Fragility of Networked Protest.

Yale University Press, New Haven, CT, USA, 2017. ISBN 0-300-21512-6 (hardcover). xxxi + 326 pp. LCCN HM742 .T84 2017.

Vacca:2010:MIS

[Vac10] John R. Vacca, editor. Managing information security. Elsevier, Amsterdam, The Netherlands, 2010. ISBN 1-59749-533-6 (papercover). xxiv + 296 pp. LCCN QA76.9.A25 M31845 2010.

Wilhelm:2011:NHU

[WA11] Thomas Wilhelm and Jason Andress. Ninja hacking: unconventional penetration testing tactics and techniques. Syngress Publishing, Inc., Rockland, MA, USA, 2011. ISBN 1-59749-588-3. xii + 310 pp. LCCN QA76.9.A25 W548 2011.

Watson:2014:SEP

[Wat14] Gavin Watson. Social engineering penetration testing: executing social engineering pen tests, assessments and defense. Syngress Publishing, Inc., Rockland, MA, USA, 2014. ISBN 0-12-420124-5. ???? pp. LCCN HM668 .W38 2014.

Weidman:2014:PTH

[Wei14] Georgia Weidman. Penetration testing: a hands-on introduction to hacking. No Starch Press, San Francisco, CA, USA, 2014. ISBN 1-59327-564-1 (paperback). ???? pp. LCCN QA76.9.A25 W4258

Wilhelm:2010:PPT

[Wil10] Thomas Wilhelm, editor. Professional penetration testing: creating and operating a formal hacking lab. Syngress Publishing, Inc., Rockland, MA, USA, 2010. ISBN 1-59749-425-9 (paperback), 1-59749-466-6 (DVD). xix + 504 pp. LCCN TK5105.59 .W544 2010.

Wilhelm:2013:PPT

[Wil13] Thomas Wilhelm, editor. Professional penetration testing. Syngress Publishing, Inc., Rockland, MA, USA, second edition, 2013. ISBN 1-59749-993-5. ???? pp. LCCN TK5105.59 .W544 2013.